the average number of automobiles per minute stopping for gas at a particular service station along the coastal road is 4 what is the probability that in any given minute more than two will stop for gas


1
Expert's answer
2022-04-29T09:12:15-0400

It has to be pointed out that the distribution is not specified. Suppose that "X" is a random variable that corresponds to the average number of automobiles per minute. The aim is to compute "P(X>2)". Consider the most typical distributions:

1. Suppose that "X" has a normal distribution :"X\sim N(\mu, \sigma^2)."

Then


"P(X>2)=1-P(X\le2)=1-P(Z\le\dfrac{2-4}{1})"

"=1-P(Z\le-2)\approx0.97725"


2. Suppose that "X" has a Poisson distribution with "\lambda=4:X\sim Po(\lambda)."

Then


"P(X>2)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dfrac{e^{-4}(4)^0}{0!}-\dfrac{e^{-4}(4)^1}{1!}"

"-\dfrac{e^{-4}(4)^2}{2!}\approx0.76190"